OpenStudy (anonymous):

I'm confused at how to do this? Use interval notation to express the solution set and graph the solution set on a number line. ((x-18)/6)>or equal to ((x-9)/9)+(1/18)

OpenStudy (anonymous):

\[\frac{x-18}{6}\geq \frac{x-9}{9}+\frac{1}{18}\] start by multiplying both sides by 18 to get \[3(x-18)\geq2(x-9)+1\]

OpenStudy (anonymous):

You have an inequality equation. Treat it as a normal equation, except when multipling or dividing both sides with a negative number, you have to "flip" the >= sign.

OpenStudy (anonymous):

Thank you both!!

OpenStudy (anonymous):

Do you know what its solution set is? Like (-37, infinity) ?

OpenStudy (anonymous):

then multiply out to get \[3x-54\geq 2x-18+1\] \[x\geq 54-18+1\] \[x\geq 37\]

OpenStudy (anonymous):

think it is \[[37,\infty)\]

OpenStudy (anonymous):

unless i made an algebra mistake

OpenStudy (anonymous):

Right, didnt mean to put the "-". Thanks!